OpenStudy (anonymous):

What value of c solves the equation? 4^c = 1/16

OpenStudy (anonymous):

@WilliamZhang

OpenStudy (anonymous):

4^c=(2^2)^c=2^(2c)=1/16=2^-4 2c=-4 c=-2